In response to complaints about high prices, a grocery chain runs the following advertising campaign: If you pay your child $5 to go buy $50 worth of groceries, then your child makes twice as much on the trip as we do. Youve collected the following information from the grocery chains financial statements ($ in millions) Sales $570.0 Net income 28.5 Total assets 470.0 Total debt 340.0 What is the profit margin for the child as a percentage of what they spend and the profit margin for the store? (Do not round intermediate calculations. Enter your answers as a percent rounded to 2 decimal places, e.g., 32.16.) What is the stores ROE? (Do not round intermediate calculations. Enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.) a. b. a. Childs profit margin Stores profit margin b. Stores ROE < Prev 3of 10m Next 25 MacBook Air
Answer #1
Child's profit margin Net income/Sales 10.00% =5/50
Store's profit margin Net income/Sales 5.00% =28.5/570
Store's ROE Net income/Total equity 21.92% =28.5/(470-340)